OpenStudy (love.kat123):

MEDAL+FAN Can you add the square root of 6 and the square root of 24? Explain why or why not.

OpenStudy (zzr0ck3r):

You can add any two numbers

OpenStudy (love.kat123):

okay

OpenStudy (zzr0ck3r):

I think what they are getting as is \(\sqrt{6}+\sqrt{24}=\sqrt{6}+\sqrt{6*4}=\sqrt{6}+\sqrt{6}\sqrt{4}=\sqrt{6}+2\sqrt{6} = 3\sqrt{6}\)

OpenStudy (zzr0ck3r):

but badly worded question indeed

OpenStudy (love.kat123):

lol okay thanks